OpenStudy (anonymous):

Use the logarithmic properties to expand ln(x^2-16/x^4), where x>4

OpenStudy (tkhunny):

Are you SURE you don't mean \(\dfrac{x^{2} - 16}{x^{4}}\)? Just checking.

OpenStudy (anonymous):

Yes, with parenthesis.

OpenStudy (tkhunny):

I'm not certain you clarified it. Which is it? \(\ln\left(x^{2} - \dfrac{16}{x^{4}}\right)\) as you have written it, or otherwise, perhaps \(\ln\left(\dfrac{x^{2} - 16}{x^{4}}\right)\)

OpenStudy (anonymous):

the second one you wrote.

OpenStudy (tkhunny):

Do you see why that is NOT what you have written? It is important that you understand this. It is a matter of the Order of Operations. Factor the numerator and see if anything dawns on you.

OpenStudy (anonymous):

Sorry, it was hard to write it what I wanted.

OpenStudy (tkhunny):

It only requires additional parentheses. I'm guessing you DO know how to do that. Take the time to write clearly. Now, how does that numerator factor?

OpenStudy (anonymous):

factor the numerator and the denominator

OpenStudy (tkhunny):

Well, yes, but factoring the denominator is trivial.

OpenStudy (inkyvoyd):

HINT: to factor the numerator observe that it is in fact a difference of squares

OpenStudy (tkhunny):

And, just for the record, factoring the Difference of Squares should be ALMOST trivial by the time you get to the calculus. So, factor away. Let's see what you get.

OpenStudy (anonymous):

and what about x>4?

OpenStudy (tkhunny):

We can talk about that later. Factor!

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

(x^4)(x^2-16)=x^6-16x^4?

OpenStudy (tkhunny):

\(\dfrac{x^{2} - 16}{x^{4}} = \dfrac{(x+4)(x-4)}{x^{4}}\) What were you doing? How did that \(x^{4}\) get into the numerator?

OpenStudy (anonymous):

I'm sorry /.\ my teacher haven't teach me yet.

OpenStudy (anonymous):

she gave me during winter break.

OpenStudy (tkhunny):

Not sure where to go with that. If you cannot factor a difference of squares, you cannot solve this problem. You'll have to wait until you cover that material. From the logarithm point of view, we have this well-known rule: log(a*b) = log(a) + log(b) This is not ALWAYS the case unless you KNOW that a and b are positive. This completely explains the restriction x>4.

OpenStudy (anonymous):

I know the log(a*b)=log(a)+log(b)

OpenStudy (tkhunny):

That's nice, but that rule is of no value whatsoever on this problem unless you can factor the numerator. Kinda' stuck where we are.

OpenStudy (anonymous):

aw, shoot. :/ Well, thanks for helping me and I am sorry I am wasting your time.

OpenStudy (tkhunny):

Nope, no time wasting at all. It is good to figure out where you are. Now we know and you know where you need a little more local help. We can't fix everything, here. Hang in there. You'll get it. :-)

OpenStudy (anonymous):

Thank you so much. :) Have a nice day.